### v1

# Emergence of generic first-passage time distributions for large Markovian networks

Abstract
First-passage times are often the most relevant aspect of a complex Markovian network, because they signify when information processing has resulted in a definite decision. Previous studies have shown that for kinetic proofreading networks in the limit of large network size the first-passage time distribution converges either to a delta or to an exponential distribution. Remarkably, these two forms correspond to the two extreme distributions of minimal and maximal entropy for a fixed mean, respectively. Here we build on the connection between first-passage times and graph theory to show that these two limits are not model-specific, but arise generically in Markovian networks from the distribution of the eigenvalues of the generator matrix. A deterministic peak emerges when infinitely many eigenvalues contribute, while the exponential limit arises from a single dominant eigenvalue. We also show that the exponential limit emerges robustly for reversible networks when a backward bias exists. In contrast, the deterministic limit is not obtained from a simple reversal of this condition, but under structurally tighter conditions, revealing a fundamental asymmetry between both regimes. Our theoretical analysis is illustrated and validated by computer simulations of one-step master equations and random networks.
preprint: ???
I Introduction
Although real world networks often show an overwhelming level of complexity [1, 2, 3], often it matters mainly when a certain absorbing state is reached, namely when this state marks the completion of a certain process [4, 5, 6, 7, 8]. This is especially true in biological systems, which often have to switch to another program once a certain state is reached. While in development this would be the completion of the organism, in cellular signaling it could be the decision to respond to an extracellular signal [9]. A prominent example of the latter case would be kinetic proofreading, a process in which the accuracy of the decision-making process is increased by energy consumption [10]. Possible outcomes include cell division in response to growth factors or activation of the immune system after detecting a pathogen [11, 12].
Biochemical processes can be described by time-continuous and space-discrete master equations [13, 14]. Reaching a certain state in such a network is called a first passage and the corresponding time is called the first-passage time (FPT) (Fig. 1(a)). Like the underlying state variable, it is a stochastic variable with a certain distribution. The statistical properties of first-passage times play a central role in physics [15, 16, 17, 18, 19, 20, 21, 22], chemistry [23, 24, 25, 26], biology [27, 8, 28, 29, 30, 31, 32, 33], and economics [34, 35, 36], where they characterize noise-driven event timing, barrier-crossing dynamics, and rare transitions in complex stochastic systems [37, 38, 39, 7].
Figure 1: Generic first-passage time (FPT) distributions for large Markovian networks. (a) FPT in a large network: a trajectory starting at an initial state (square) reaches the target state NN (diamond) for the first time. (b) Two generic distributions have been observed before to emerge for kinetic proofreading networks in the limit of very large systems: the delta and the exponential distributions. (c) Here we show that in the limit $N\to\infty$, the FPT-distribution is determined by the eigenvalue structure of the generator matrix. If infinitely many eigenvalues contribute, the distribution collapses to a delta peak. If $\lambda_{1}^{-1}$ dominates the sum of inverse eigenvalues, the FPT rescaled by its mean converges to an exponential law. (d) Examples of graph theoretical concepts used here: a graph, a spanning tree and a two-tree spanning forest. A graph-theoretical decomposition provides an exact representation of the FPT-moments and the Laplace transform of the FPT-density.
In the context of master equation models for kinetic proofreading, it has been observed before, both through analytical and numerical means, that on the timescale of the mean completion time, the FPT density tends to converge to one of two simple forms in the limit of large network size: either to a delta distribution, indicating quasi-deterministic behavior, or to an exponential distribution, indicating quasi-memoryless behavior [40, 41] (Fig. 1(b)). As pointed out by the authors, this phenomenon helps to explain the success of coarse-grained Markov chain models for complex systems. It also raises the question how much microscopic information can be inferred from macroscopic observations. In fact, it is common practice to infer system properties from experimental first-passage time measurements [42, 43, 44, 45, 46, 47, 48]. If only the delta distribution or the exponential distribution, both of which are fully characterized by their means, emerged in the limit of large networks, such inference methods would become impossible in the limit of large systems. In the study of complex networks, ignorance about the microscopic details of the network remains a substantial obstacle to understanding these systems [49, 50, 51]. Model independent results are therefore particularly desirable [52]. In this context, Godec and Metzler [53, 54] identified the separation of time scales between direct and indirect trajectories as a mechanism leading to universal FPT-asymptotics.
Interestingly, the two limiting distributions identified for kinetic proofreading networks also figure prominently in information theory and stochastic thermodynamics, because they correspond to the two extreme regimes of entropy for the first-passage time distribution: while the delta distribution is deterministic and contains no uncertainty, the exponential distribution can be shown to correspond to the case of maximal entropy [55, 56, 57, 58]. Thus both distributions make it difficult to infer the underlying microscopic dynamics, but they must emerge for very different reasons.
Here we show how the two emerging limit distributions identified earlier for kinetic proofreading are in general related to the distribution of the eigenvalues of the generator matrix. We show that the key quantity is the sum of the inverse eigenvalues of the generator matrix. If infinitely many eigenvalues contribute comparably, a delta function emerges, while the exponential function results, when one value $\lambda_{1}^{-1}$ dominates (Fig. 1(c)). However, these two cases are not simple mirrors of each other, because the delta distribution requires stricter conditions than the exponential distribution. The deep connection between eigenspectra and properties of the FPT distribution - especially in the case where detailed balance holds for the steady state distribution, ensuring that the spectrum is real - has proven fruitful in the context of extreme value and time-average statistics [59, 60]. Here we analyze this connection using the interpretation of first-passage times of Markovian networks in terms of graph theory, which recently gained more interest in the context of first passage time problems [61, 62, 63, 64, 65, 66, 67, 68, 69]. In particular, we will use the fact that first passage time moments can be obtained from a decomposition of the Markovian network into spanning trees and two-tree spanning forests (Fig. 1(d)). We illustrate and validate our theoretical analysis by stochastic computer simulations using the Gillespie algorithm.
This paper is structured as follows. Sec. 2 gives an overview on FPT-theory for Markovian networks, including the graph theoretical interpretation and general results on the spectrum of the generator matrix. In Sec. III.2, we consider the limit of large networks by relating the dependence of the cumulants of the FPT to the distribution of the eigenvalues depending on the limiting ratio of the weights of the two-tree spanning forests of the network. Then, we give sufficient conditions to conclude convergence to the deterministic and exponential limits in Secs. III.3 and III.4, respectively. Moreover, we demonstrate in Sec. III.5 that the naive expectation that a forward-bias would lead to the deterministic limit is misleading. Finally, we discuss this and other caveats to the universality of FPT distributions for large networks in Sec. IV, uncovering which systems will lead to simple, large-scale statistics and which ones will not. Computer simulations with the Gillespie algorithm will be used throughout to illustrate and validate our theoretical results.
II First-Passage Time Theory for Markovian Networks
Figure 2: The first-passage time (FPT) of a one-step master equation compared to its mean as system size NN increases. (a) The network of the one-step master equation with site-dependent forward rates $k_{i}$ and backward rates $r_{i}$. (b) FPT-statistics for $10^{5}$ runs with randomly drawn rates. Forward bias: rates drawn from uniform distributions, $k_{i}\sim U[1,2]$, $r_{i}\sim U[0,1]$, showing convergence to a delta-distribution (deterministic limit). Backward bias: rates drawn from uniform distributions, $k_{i}\sim U[0.5,1]$, $r_{i}\sim U[0.5,1.5]$, showing convergence to an exponential distribution (exponential limit). (c) FPT-statistics for $10^{5}$ runs with constant rates. Forward bias: rates chosen as $k_{i}=\frac{3}{2}$, $r_{i}=\frac{1}{2}$, showing convergence to a delta-distribution (deterministic limit). Backward bias: rates chosen as $k_{i}=\frac{3}{4}$, $r=1$, showing convergence to an exponential distribution (exponential limit).
II.1 Master equation
Consider a time-homogeneous jump process on $N$ states governed by the master equation
$\dot{p}_{i}(t)=\sum_{j=1}^{N}k_{ji}p_{j}(t)-k_{ij}p_{i}(t),$ (1)
with $p_{i}(t)$ the probability of being in state $i$ and $k_{ij}$ the jump rate from $i$ to $j$. Defining
$\hat{K}_{ij}=\begin{cases}k_{ji},&i\neq j,\\ -\sum_{m}k_{im},&i=j,\end{cases}$ (2)
the equation becomes $\dot{\mathbf{p}}(t)=\hat{K}\mathbf{p}(t)$, with $\sum_{i=1}^{N}p_{i}(t)=1$ and $\hat{K}$ singular. The discussion can be restricted to irreducible Markov chains, so that $\hat{K}$ is guaranteed to have a simple zero eigenvalue, and all other eigenvalues have a negative real part. The corresponding eigenvector is the steady-state distribution $\mathbf{p}_{s}$. One calls the process reversible, if $\mathbf{p}_{s}$ obeys the detailed balance condition, $k_{ij}p_{s,i}=k_{ji}p_{s,j}$ $\forall i,j$. Then $\hat{K}$ is similar to a symmetric matrix $\hat{K}^{s}=P\hat{K}P^{-1}$, where $P=\text{diag}(\sqrt{p_{s,1}},...,\sqrt{p_{s,N}})$, which is indeed symmetric, since for $i\neq j$:
$\hat{K}^{s}_{ij}$ $=k_{ij}\sqrt{\frac{p_{i}}{p_{j}}}=k_{ji}\sqrt{\frac{p_{j}}{p_{i}}}=\hat{K}^{s}_{ij}.$ (3)
In particular, then all eigenvalues of $\hat{K}$ are real.
II.2 Example: one-step master equation
As an illustrative example, we consider a one-step master equation or birth-death process:
$\dot{p}_{i}=-(k_{i}+r_{i})p_{i}+k_{i-1}p_{i-1}+r_{i+1}p_{i+1},$ (4)
with $0\leq i\leq N$, a reflecting boundary at $i=0$ and finite rates ($k_{\text{min}}\leq k_{i}\leq k_{\text{max}}$, $r_{\text{min}}\leq r_{i}\leq r_{\text{max}}$ for all $i$). The steady state obeys the detailed balance condition, $k_{i}p_{s,i}=r_{i+1}p_{s,i+1}$ for all $i$ due to the reflecting boundary. The corresponding network is shown in Fig. 2(a).
It is well known that the mean first-passage time (MFPT) to reach state $N$ starting from state $m$ can be expressed as
$\langle\tau\rangle_{m\to N}=\sum_{i=m}^{N-1}\sum_{j=1}^{i}\frac{1}{k_{j}}\prod_{l=j+1}^{i}\frac{r_{l}}{k_{l}}.$ (5)
This result is typically derived using recursive arguments [39, 70], but can also be obtained directly within the graph-theoretical framework [62, 67]. In contrast, determining the full FPT time distribution is substantially more challenging. Exact solutions exist only for particular choices for the rates [6, 71], while most available results rely on approximations [72, 73, 74, 75, 76, 77].
As an alternative to analytical calculations, FPT-distributions can be efficiently sampled using stochastic simulations based on the Gillespie algorithm [78, 79]. In Fig. 2(b), we show the results of such simulations for the one-step master equation from Eq. 4. We first use random rates drawn from uniform distributions $k_{i}\sim U[1,2]$, $r_{i}\sim U[0,1]$, so that the process has a forward bias. System size $N$ is increased from $10$ through $100$ to $1000$. One nicely sees how the FPT distribution becomes sharply peaked, approaching a delta-like form on the scale of its mean. In marked contrast, if the rates are drawn from $k_{i}\sim U[0.5,1]$, $r_{i}\sim U[0.5,1.5]$, we have a backward bias and the distribution converges toward an exponential form.
Fig. 2(c) shows the same FPT statistics for constant rates instead of randomly chosen ones, with $(k_{i},r_{i})=(\frac{3}{2},\frac{1}{2})$ to get a forward bias and $(k_{i},r_{i})=(\frac{3}{4},1)$ to get a backward bias. The limiting behavior of the distributions is qualitatively identical to the previous case of random rates, showing that the limiting behaviors are not a consequence of randomly assigning the rates. This confirms earlier results for kinetic proofreading networks that these two distributions tend to emerge in the limits of large system sizes, and that they can have a relation to the bias in the system [40, 41].
II.3 First-passage time and graph theory
For the first-passage time (FPT) to state $N$, one can restrict the system to the first N1N-1 components:
$\dot{\mathbf{p}}(t)=K\mathbf{p}(t),$ (6)
where $K$ is the $N$-th principal minor of $\hat{K}$. The FPT density obeys the adjoint master equation [38, 39]:
$\dot{\mathbf{f}}(t)=K^{T}\mathbf{f}(t)\quad\Rightarrow\quad\mathbf{f}(t)=e^{K^{T}t}\mathbf{f}_{0},$ (7)
with $f_{0,i}=k_{i,N}$. This implies that $\mathbf{f}(t)$ belongs to the class of phase-type distributions [80]. Its Laplace transform reads:
$\mathcal{L}{\mathbf{f}}(s)=(K^{T}-sI)^{-1}\mathbf{f}_{0}=\sum_{n=0}^{\infty}s^{n}(K^{T-1})^{n}\mathbf{1},$ (8)
so the $n$-th FPT moment is given by:
$\boldsymbol{\tau}^{(n)}=n(-K^{T})^{-1}\boldsymbol{\tau}^{(n-1)}=n!\,((-K^{T})^{-1})^{n}\mathbf{1}.$ (9)
Figure 3: Illustration of graph theoretical concepts. (a) Example of a Markovian network. (b) A spanning tree of the graph with root in 55. (c) A two-tree spanning forest with roots in $3$ and $5$. (d) Network with an outgoing edge to 55 with weight $s$ added to all vertices, which can be used to calculate the Laplace transform of the first-passage time.
For notational clarity, we write $\tau^{(1)}_{i}=\langle\tau_{m}\rangle$ for the mean first-passage time (MFPT) and define $M:=(-K^{T})^{-1}$ in the following. The components of this matrix have a graph-theoretical interpretation originating from the All-Minors Matrix-Tree Theorem, a generalization of Kirchhoff’s theorem (see also Supplement I) [81, 61, 63]:
$M_{ij}=\frac{\sum_{\mathcal{F}^{i\to j}_{[j,N]}}w(\mathcal{F})}{\sum_{\mathcal{T}_{[N]}}w(\mathcal{T})}$ (10)
where $\mathcal{T}_{[N]}$ denotes the sum over all spanning trees rooted at $N$ and $\mathcal{F}^{i\to j}_{[j,N]}$ are the two-tree spanning forests with one tree rooted at $N$ and the other one rooted at $j$ containing $i$. The weight $w(G)$ of graph $G$ is defined as the product over all the edges (rates) of $G$, i.e., $w(G):=\prod_{k_{ij}\in E(G)}k_{ij}$. In particular, this implies that $M_{ij}$ is non-negative $\forall i,j$ and all components are positive if and only if the graph is strongly connected (ergodic) except for maybe $N$ (there is a path from any vertex excluding $N$ to any other vertex). Applying this identity to Eq. (9), one obtains, for instance, for the mean:
$\langle\tau_{m}\rangle=\frac{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{m\to j}_{[j,N]}}w(\mathcal{F})}{\sum_{\mathcal{T}_{[N]}}w(\mathcal{T})},$ (11)
and one can also get similar expressions for higher moments. In Fig. 3(a), we show as an example a Markovian network with five vertices. Fig. 3(b) and (c) show examples for a spanning tree and for a two-tree spanning forest of that network, respectively.
An immediate corollary from the graph-theoretic interpretation is that $M_{ij}\leq M_{jj}$ $\forall i,j$, since every tree rooted at $j$ containing $i$ is also a tree rooted at $j$ containing $j$. Hence,
$\max_{i}\langle\tau\rangle_{i}$ $=\max_{i}\Big(\sum_{j}M_{ij}\Big)\leq\text{tr}(M)=\sum_{i}\frac{1}{\lambda_{i}}.$ (12)
We next observe that $s\mathds{1}-K^{T}$ is of the same shape as $K^{T}$ if one replaces $k_{i,N}\to k_{i,N}+s$, as illustrated in Fig. 3(d). So $s\mathds{1}-K^{T}$ can be inverted by the same logic:
$(s\mathds{1}-K^{T})^{-1}_{ij}=\frac{\sum_{\mathcal{F}_{s,[j,N]}^{i\rightarrow j}}w(\mathcal{F})}{\sum_{\mathcal{T}_{s,[N]}}w(\mathcal{T})}.$ (13)
The index $s$ indicates that here the $s$ rates have to be included. In particular, note that the denominator is the characteristic polynomial of $K^{T}$:
$\sum_{\mathcal{T}_{s,[N]}}w(\mathcal{T})=\det(s\mathds{1}-K^{T})=\prod_{i=1}^{N-1}(s+\lambda_{i}),$ (14)
where $\text{Re}(\lambda_{1})\leq\text{Re}(\lambda_{2})\leq...\leq\text{Re}(\lambda_{N-1})$ are the eigenvalues of $K^{T}$
For the FPT density, one then finds:
$\mathcal{L}{f}_{m}(s)$ $=\frac{\sum_{j=1}^{N-1}k_{j,N}\sum_{\mathcal{F}_{s,[j,N]}^{m\rightarrow j}}w(\mathcal{F})}{\sum_{\mathcal{T}_{s,[N]}}w(\mathcal{T})}$ (15)
$=\frac{\sum_{j=1}^{N-1}k_{j,N}\sum_{\mathcal{F}_{s,[j,N]}^{m\rightarrow j}}w(\mathcal{F})}{\sum_{n=0}^{N-1}\Big(\sum_{\begin{subarray}{c}|\mathcal{F}_{[N]}|=N-1-n\end{subarray}}w(\mathcal{F})\Big)s^{n}},$ (16)
where the second sum in the denominator denotes the sum over all spanning forests with one root in $N$ and N1nN-1-n edges (meaning that it is an $n$-tree forest).
III Limit of large Networks
III.1 Perron-Frobenius theorem and an intuitive argument
The Perron-Frobenius theorem guarantees the existence of a simple, dominant real eigenvalue, which is known as the Perron-Frobenius eigenvalue, for positive and, more generally, for non-negative matrices (with strict dominance in the positive case) [82]. This theorem is a cornerstone in the study of Markov chains, as the existence of a steady-state solution, and its uniqueness for irreducible chains, follow directly from it [83, 84].
As argued above, $M$ is positive if the network is strongly connected, so $\frac{1}{\lambda_{1}}$ corresponds to the Perron–Frobenius eigenvalue. Furthermore, it is bounded by the smallest and largest row sums of $M$ [85]:
$\min_{i}\langle\tau_{i}\rangle\leq\frac{1}{\lambda_{1}}\leq\max_{i}\langle\tau_{i}\rangle.$ (17)
Combining this with Eq. (12) yields upper and lower bounds for the maximum MFPT:
$\frac{1}{\lambda_{1}}\leq\max_{i}\langle\tau_{i}\rangle\leq\sum_{i}\frac{1}{\lambda_{i}}.$ (18)
If, in the limit $N\to\infty$, the leading term dominates such that
$\lim_{N\to\infty}\frac{\frac{1}{\lambda_{1}}}{\sum_{i}\frac{1}{\lambda_{i}}}=1,$ (19)
then the MFPT scales as $\langle\tau\rangle:=\max_{i}\langle\tau_{i}\rangle\sim\frac{1}{\lambda_{1}}$, and one expects $\frac{1}{\langle\tau\rangle}M\sim P$, where $P$ is the projection operator onto the Perron-Frobenius eigenspace. Using Eq. (9), this implies for the moments:
$\boldsymbol{\tau}^{(n)}\sim n!P^{n}\mathbf{1}=n!P\mathbf{1}=n!\boldsymbol{\tau}^{(1)},$ (20)
since the projection operator satisfies $P^{2}=P$. Thus, one obtains exactly the moments of an exponential distribution.
Conversely, if $\frac{1}{\lambda_{1}}\ll\langle\tau\rangle$, one expects
$\langle\tau\rangle\sim\sum_{i}\frac{1}{\lambda_{i}},\qquad\sigma_{\tau}^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}\sim\sum_{i}\frac{1}{\lambda_{i}^{2}},$ (21)
which leads to:
$\frac{\sigma_{\tau}^{2}}{\langle\tau\rangle^{2}}\sim\frac{\sum_{i}\frac{1}{\lambda_{i}^{2}}}{\left(\sum_{i}\frac{1}{\lambda_{i}}\right)^{2}}\overset{N\to\infty}{\longrightarrow}0,$ (22)
provided that infinitely many terms $\frac{1}{\lambda_{i}}$ contribute non-negligibly to the sum. In this case, the first-passage process approaches a deterministic limit as $N\to\infty$.
While this reasoning is heuristic - for instance, it ignores possible changes in the Perron-Frobenius eigenvector as $N\to\infty$ - it nevertheless provides valuable intuition for the analysis that follows, where we establish under which conditions this argument is valid and discuss cases in which it fails.
III.2 Macroscopic forest condition
Intuitively, the large-network limit can yield meaningful macroscopic behavior only if the initial state $m$ is sufficiently far from the target state $N$. If the system starts too close to $N$, a small, local subnetwork may dominate the first-passage dynamics, leading to non-universal asymptotic behavior characteristics of the full system. As argued in the previous section, the deterministic and exponential limits depend on the distribution of the eigenvalues. We now translate this insight into a mathematically precise condition.
First, observe that
$\frac{1}{1+\frac{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{m\to j}_{[j,N]}}w(\mathcal{F})}{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{m\to N}_{[j,N]}}w(\mathcal{F})}}=1-\frac{\langle\tau_{m}\rangle}{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}}},$ (23)
which follows from Eq. (11) and from the identity $\mathrm{tr}(M)=\sum_{j=1}^{N-1}\sum_{\mathcal{F}_{[j,N]}}w(\mathcal{F})$. Thus, one concludes that
$r:=\lim_{N\to\infty}\frac{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{m\to N}_{[j,N]}}w(\mathcal{F})}{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{m\to j}_{[j,N]}}w(\mathcal{F})}<\infty$ (24)
$\Leftrightarrow$ $\lim_{N\to\infty}\frac{\langle\tau_{m}\rangle}{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}}}=\frac{1}{1+r}>0.$ (25)
Therefore, if the ratio $r$ of the two-tree spanning forests in which the vertex $m$ lies on the $N$-tree to those in which it lies on the complementary tree remains finite as $N\to\infty$, then the mean first-passage time (MFPT) from $m$ to $N$ scales with the sum of the eigenvalue inverses of $M$. Note that this condition reflects the requirement that $m$ must not be located too close to $N$, since otherwise $m$ would almost always belong to the $N$-tree, violating the macroscopic scaling assumption.
An upper bound on $r$ with a more intuitive interpretation is given by
$r\leq\max_{j=1,\ldots,N-1}\frac{p_{m\to\{j,N\}}(N)}{1-p_{m\to\{j,N\}}(N)},$ (26)
where $p_{m\to\{j,N\}}(N)$ denotes the splitting probability that a process starting at $m$ reaches $N$ before reaching the other vertex $j$. This bound implies that $r$ remains finite in the limit $N\to\infty$, unless there exists another vertex whose probability of being reached before vertex $N$ vanishes in that limit.
Using Eq. (15) to express the Laplace transform of the FPT density with time normalized to the mean, one finds for the logarithm of $\mathcal{L}f(s)$:
$\ln(\mathcal{L}{f}(s))=\sum_{n=1}^{\infty}\frac{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}^{n}}}{\langle\tau_{m}\rangle^{n}}\frac{(-s)^{n}}{n}+\ln\sum_{n=0}^{N-1}r_{n}s^{n},$ (27)
where:
$r_{n}:=\frac{\sum_{i_{1},...,i_{n}}\sum_{\mathcal{F}_{[i_{1},...,i_{n},N]}^{m\rightarrow N}}w(\mathcal{F})}{\langle\tau_{m}\rangle^{n}\sum_{\mathcal{T}_{[N]}}w(\mathcal{T})}.$ (28)
Observe that Eq. (11) implies that $r_{1}=r$. Exploiting the graph theoretical interpretation of the coefficients, one can show that for any $N$ holds (see Supplement II):
$1\leq$ $\sum_{n=0}^{N-1}r_{n}s^{n}\leq\frac{1-s}{1-(1+r)s},$ (29)
for $s<\frac{1}{1+r}$. In particular, if $r=0$, this implies:
$\ln(\mathcal{L}{f}(s))=\sum_{n=1}^{\infty}\lim_{N\to\infty}\frac{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}^{n}}}{\langle\tau_{m}\rangle^{n}}\frac{(-s)^{n}}{n},$ (30)
and one concludes that the $n$-th cumulant is given by:
$\lim_{N\to\infty}\kappa_{n}\Big(\frac{t}{\langle\tau_{m}\rangle}\Big)=(n-1)!\lim_{N\to\infty}\frac{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}^{n}}}{\Big(\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}}\Big)^{n}}.$ (31)
This result is the key insight that allows us to understand the two limiting behaviors: if there exists a single dominant eigenvalue of $K^{-1}$ in the limit $N\to\infty$ such that $\frac{\frac{1}{\lambda_{1}}}{\sum_{i}\frac{1}{\lambda_{i}}}\to 1$, then the cumulants become $\kappa_{n}\to(n-1)!$, so the FPT distribution converges to an exponential distribution. Conversely, if infinitely many eigenvalues remain of comparable magnitude, the cumulants become $\kappa_{n}\to\delta_{1n}$. So the distribution collapses to a delta peak, corresponding to the deterministic limit. In the following, we analyze these two regimes separately and derive explicit conditions on the eigenvalue spectrum under which each limiting behavior arises.
Note that if there is a unique path from $m$ to $N$ that contains all vertices (i.e., it is a Hamiltonian path), then there cannot be any spanning forest of more than one tree with a $m\to N$ path. Hence, the macroscopic network condition is true and yields $r=0$ even before taking the limit $N\to\infty$. This is the case in the example of the one step master equation from Sec. II.2 when choosing $m=1$ and also for the kinetic proof-reading networks [40, 41].
III.3 Deterministic limit
For the deterministic limit, it is necessary and sufficient that the variance $\kappa_{2}\big(\frac{t}{\langle\tau_{m}\rangle}\big)=\sigma^{2}_{m}$ vanishes as $N\to\infty$. If the macroscopic forest condition holds with $r=0$, this requires:
$\frac{\sum_{i=1}^{N}\frac{1}{\lambda_{i}^{2}}}{\Big(\sum_{i=1}^{N}\frac{1}{\lambda_{i}}\Big)^{2}}\overset{N\to\infty}{\longrightarrow}0.$ (32)
In Supplement III, we prove that this holds if $\langle\tau_{m}\rangle\overset{N\to\infty}{\longrightarrow}\infty$ and $\lim_{N\to\infty}\lambda_{1}\geq M>0$. The former condition that the MFPT from $m$ to $N$ diverge simply reflects the fact again that one has to start macroscopically far away from the target $N$ in order for the large network limit to give any generic results. The latter condition requires that the smallest eigenvalue $\lambda_{1}$, which is real as argued before, stay bounded away from 0 as $N\to\infty$.
If the network is strongly connected, that is, any two vertices are connected by a path, one can use a result by Lawler and Sokal [86] who generalized a famous result by Cheeger [87]. If the graph of $\hat{K}$ is strongly connected, then holds for the eigenvalue of $\hat{K}$ with the second smallest real part $\mu_{2}$:
$\frac{\Phi^{2}}{8c}\leq\text{Re}(\mu_{2}),$ (33)
where $c:=\max_{i}\sum_{j}k_{ij}$ is the maximal sum of the outgoing rates from a vertex and the conductance $\Phi$ is defined as:
$\Phi:=\min_{0<p_{s}(A)<1}\frac{\sum_{i\in A,j\in A^{c}}k_{ij}p_{s,i}}{p_{s}(A)p_{s}(A^{c})},$ (34)
where $p_{s,i}$ is the steady state distribution of the network and the minimum is taken over all partitions of the vertex set into $A$ and $A^{c}$. Note that, if the detailed balance condition holds, $A$ is a set minimizing $\Phi$ if and only if $A^{c}$ is a set minimizing $\Phi$.
This is a statement on the eigenvalues of the unreduced generator matrix $\hat{K}$. However, relevant for our case are the eigenvalues of the $N$-th principal minor $K$, which only coincide with the non-zero eigenvalues of $\hat{K}$ if $N$ is chosen to be absorbing, meaning that the network cannot be strongly connected anymore. To reconcile this, one can introduce a small back-rate $\epsilon$ from N to any other vertex $i$ and let $\epsilon\to 0$. Note that K is invariant under any choice of rates leaving $N$ as this cannot change the FPT behavior to reach $N$.
We can apply this result to the one-step master equation from Sec. II.2 starting at $m=1$. Here, one sees that a minimizing set has to be of the shape $A=\{1,...,n\}$, which yields:
$\frac{\sum_{i\in A,j\in A^{c}}k_{ij}p_{s,i}}{p_{s}(A)p_{s}(A^{c})}\geq\frac{1}{\sum_{i=1}^{n}\frac{1}{k_{i}}\prod_{j=i+1}^{n}\frac{r_{i}}{k_{i}}}=\langle\tau\rangle_{n-1\to n}^{-1}$ (35)
So, if the average time to advance one step is always finite, one can conclude convergence to the deterministic limit, explaining the convergence observed in Fig. (2)(b). Note that such a local forward bias is required to obtain the deterministic limit. A global forward bias is not sufficient, as illustrated in Sec. III.5. However, the intuition that stochastic dynamics biased to the target can be turned into a mathematically rigorous condition by considering the mean residual life (MRL) of the process [88, 89]:
$\langle\tau\rangle(t)=\frac{\sum_{i=1}^{N-1}\langle\tau_{i}\rangle p_{i}(t)}{1-p_{N}(t)},$ (36)
which describes the MFPT over time given that the process has not reached state $N$ yet. As a direct consequence of Eq. (9), one sees that the MRL obeys the following differential equation:
$-1=\frac{\text{d}}{\text{d}t}\langle\tau\rangle(t)-\frac{\dot{p}_{N}(t)}{1-p_{N}(t)}\langle\tau\rangle(t).$ (37)
The stationary solution for $t\to\infty$ is therefore given by:
$\lim_{t\to\infty}\langle\tau\rangle(t)$ $=\lim_{t\to\infty}\frac{1-p_{N}(t)}{\dot{p}_{N}(t)}=-\lim_{t\to\infty}\frac{f(t)}{\dot{f}(t)}=\frac{1}{\lambda_{1}},$ (38)
using l’Hôpital’s rule and the fact that $f(t)=\dot{p}_{N}(t)$ for choosing vertex $N$ as absorbing. The last equality is a result of the asymptotic scaling of the first-passage time density as $f(t)\sim e^{-\lambda_{1}t}$ [54]. Hence, if the MRL becomes small for long times relative to the MFPT (which is the MRL at $t=0$), this allows to conclude convergence to the deterministic limit. In the case that the network is not strongly connected, the FPT-distributions for the connected components have to be computed. The Laplace transform of the global FPT can then be expressed as:
$\mathcal{L}{f}(s)=\sum_{i}P_{i}\prod_{j=1}^{n_{i}}\mathcal{L}{g}_{ij}(s),$ (39)
where $P_{i}$ is the probability to take a specific path through the connected components and $\mathcal{L}{g}_{ij}(s)$ are FPTs of the connected components, hence satisfying $\mathcal{L}{g}_{ij}(0)=1$, $\frac{\text{d}\mathcal{L}{g}_{ij}}{\text{d}s}\Big|_{s=0}=-\langle\tau_{ij}\rangle$. Then $\langle\tau\rangle_{i}=\sum_{j}\langle\tau\rangle_{ij}$ is the MFPT along the $i$-th path so that $\langle\tau\rangle=\sum_{i}P_{i}\langle\tau\rangle_{i}$. If $\frac{\langle\tau\rangle_{ij}}{\langle\tau\rangle}\to 0$ $\forall j$ as $N\to\infty$, one gets for $s\to\frac{s}{\langle\tau\rangle}$:
$\ln\Big(\prod_{j=1}^{n_{i}}\mathcal{L}{g}_{ij}\Big(\frac{s}{\langle\tau\rangle}\Big)\Big)=-\frac{\langle\tau\rangle_{i}}{\langle\tau\rangle}s+\mathcal{O}\Big(\Big(\frac{s}{\langle\tau\rangle}\Big)^{2}\Big),$ (40)
so, as $N\to\infty$, one gets:
$\mathcal{L}{f}(s)=\sum_{i}P_{i}e^{-\frac{\langle\tau\rangle_{i}}{\sum_{i}P_{i}\langle\tau\rangle_{i}}s}$ (41)
$\Rightarrow$ $f(t)=\sum_{i}P_{i}\delta\big(t{-\frac{\langle\tau\rangle_{i}}{\sum_{i}P_{i}\langle\tau\rangle_{i}}}\big),$ (42)
so the FPT-density becomes a sum over delta functions weighted by the probabilities for the corresponding paths.
As a minimal example, consider the network shown in Fig. (4). It consists of of two Poisson processes with rates $k_{1}$ and $k_{2}$, both going from vertex $1$ to vertex $N$. Since all transitions are irreversible, every vertex on its own is a connected component, so the global MFPT clearly dominates the individual steps. There are two possible paths with probabilities $P_{1}=\frac{k_{1}}{k_{1}+k_{2}}$ and $P_{2}=\frac{k_{2}}{k_{1}+k_{2}}$ with individual MFPTs $\langle\tau\rangle_{1}=\frac{N-1}{2k_{1}}$ and $\langle\tau\rangle_{2}=\frac{N-1}{2k_{2}}$, respectively. Thus, the MFPT is $\langle\tau\rangle=\frac{N-1}{k_{1}+k_{2}}$, and according to Eq. (42), the FPT density in the limit of $N\to\infty$ should read:
$f(t)$ $=\frac{k_{1}}{k_{1}+k_{2}}\delta\Big(t-\frac{k_{1}+k_{2}}{2k_{1}}\Big)+\frac{k_{2}}{k_{1}+k_{2}}\delta\Big(t-\frac{k_{1}+k_{2}}{2k_{2}}\Big).$ (43)
This result agrees well with intuition and also with the limit of the simulated first-passage time distributions shown in Fig. 4(b), which in this case is simply a weighted sum of two Erlang distributions.
Figure 4: A minimal example to illustrate how irreversible transitions can result in two delta peaks. (a) A network branching into two Poisson processes with rates $k_{1}$ and $k_{2}$, both starting at vertex $1$ and ending in vertex $N$. (b) The first-passage time density results in two distinct delta peaks as NN increases, corresponding to the two chains. Here shown for $k_{1}=1$ and $k_{2}=2$.
III.4 Exponential limit
To establish convergence to the exponential limit, one must show that
$\frac{\frac{1}{\lambda_{1}}}{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}}}\overset{N\to\infty}{\longrightarrow}1.$ (44)
Once this condition is met, the bounds on $\max_{i}\langle\tau_{i}\rangle$ (Eq. (18)) imply that
$\frac{\max_{i}\langle\tau_{i}\rangle}{\sum_{i=1}^{N-1}\frac{1}{\lambda_{i}}}\overset{N\to\infty}{\longrightarrow}1.$ (45)
Thus, the macroscopic forest condition automatically holds with $r=0$ for $\langle\tau_{m}\rangle=\max_{i}\langle\tau_{i}\rangle$. However, proving the existence of such a dominant eigenvalue is generally nontrivial. In the following, we introduce two possible tests for this limit.
The first approach is to verify that the macroscopic forest condition holds with $r=0$. Unlike in the deterministic limit, convergence to an exponential distribution requires that all cumulants approach those of the exponential distribution for all orders $n$. The requirement that the coefficient of variation converges to unity,
$\lim_{N\to\infty}\frac{\sigma_{m}^{2}}{\langle\tau_{m}\rangle^{2}}=1,$ (46)
is therefore necessary but not sufficient. Although a coefficient of variation equal to $1$ is indicative of exponential form, it could also arise in other distributions.
However, if the macroscopic forest condition holds with $r=0$, this criterion becomes sufficient, as shown in Supplement IV. This result also justifies earlier work on kinetic proofreading networks [40, 41], which inferred exponential convergence from observing that the coefficient of variation tends to $1$.
A second test is based on the physical expectation proposed earlier that the first-passage time distribution approaches a delta or exponential form depending on whether the dynamics is biased toward or away from the target state, respectively [40, 41]. If the steady-state distribution satisfies detailed balance, this expectation can be made precise. Specifically, for $\langle\tau\rangle_{m\to N}:=\max_{i}\langle\tau\rangle_{i\to N}$, one has
$\frac{\langle\tau\rangle_{N\to m}}{\langle\tau\rangle_{m\to N}}\overset{N\to\infty}{\longrightarrow}0\Rightarrow\lim_{N\to\infty}\kappa_{n}\!\left(\frac{t}{\langle\tau_{m}\rangle}\right)=(n-1)!,$ (47)
provided that the macroscopic forest condition holds for the network with $m$ as the target and $N$ as the initial state with any finite $r$:
$\lim_{N\to\infty}\frac{\sum_{j=1}^{N-1}\sum_{\mathcal{F}^{N\to m}_{[j,m]}}w(\mathcal{F})}{\sum_{j=1,j\neq m}^{N}\sum_{\mathcal{F}^{N\to j}_{[j,m]}}w(\mathcal{F})}=r<\infty.$ (48)
In this case, the first-passage time distribution for $m\to N$ converges to an exponential form. This follows from the interlacing property of the eigenvalues of different principal minors of $M$ when detailed balance holds [90]. A formal proof is provided in the Supplement IV, where we exploit that $K$ is similar to a symmetric matrix via the diagonal similarity transformation $P=\text{diag}(\sqrt{p_{s,1}},\ldots,\sqrt{p_{s,N}})$, as expained above Sec. (II.1). The argument closely follows Fisk’s short proof of Cauchy’s interlacing theorem [91].
This result applies directly to the one-step master equation. As explained in Sec. II.2, its steady-state probability distribution satisfies the detailed balance condition. Moreover, the reverse transition $N\!\to\!1$ forms a Hamiltonian path, so the macroscopic forest condition also holds even with $r=0$. This rigorously establishes the exponential limit observed in Fig. 2(c).
III.5 Global forward bias not sufficient for deterministic limit
Figure 5: FPT-statistics for a one-step master equation with $N=50$ for $10^{5}$ trials with rates as specified in Eq. (49). Despite an apparent forward bias, the limiting behavior of the distribution is exponential, showing that the expectation that a forward bias would always lead to a quasi-deterministic behavior is not correct.
Motivated by the result in the exponential regime, one might expect that the condition $\frac{\langle\tau\rangle_{1\to N}}{\langle\tau\rangle_{N\to 1}}\overset{N\to\infty}{\longrightarrow}0$ could ensure deterministic behavior. The following counterexample demonstrates that this is not the case: consider the one-step master equation from Eq. (4) with the rates
[…]
Inserting these into Eq. (5), one has:
$\langle\tau\rangle_{1\to N}$ $=\Big(\frac{3}{2}\Big)^{\frac{N}{2}+1}+\mathcal{O}(N)$ (50)
$\langle\tau\rangle_{N\to 1}$ $=3^{\frac{N}{2}+1}+\mathcal{O}(N).$ (51)
Therefore, $\lim_{N\to\infty}\frac{\langle\tau\rangle_{1\to N}}{\langle\tau\rangle_{N\to 1}}=0$. Despite this apparent forward bias, the FPT from state $1$ to state $N$ is dominated by the segment $\frac{N}{2}\to N$, where the system experiences a backward bias. This segment alone leads to an exponential limit, preventing convergence to the deterministic limit (see Fig. 5).
We can further see the inadequacy of bias-based criteria by examining the harmonic mean of the local ratios $\frac{k_{i}}{r_{i}}$:
$\frac{N-2}{\sum_{i=2}^{N-1}\frac{r_{i}}{k_{i}}}=\frac{N-2}{\frac{N-2}{2}(\frac{1}{3}+\frac{3}{2})}=\frac{12}{11}>1,$ (52)
Since the harmonic mean is always less than or equal to the geometric mean, which in turn is less than or equal to the arithmetic mean, we also have:
$1<\Big(\prod_{i=2}^{N-1}\frac{k_{i}}{r_{i}}\Big)^{\frac{1}{N-2}}\leq\frac{1}{N-2}\sum_{i=2}^{N-1}\frac{k_{i}}{r_{i}}.$ (53)
Thus, defining a ’forward bias’ through any mean of the ratios $\frac{k_{i}}{r_{i}}$ still fails to guarantee convergence to the deterministic. Moreover, unbounded rates allow for transition rates that are macroscopically large, preventing FPT universality (see Supplement V).
III.6 First-passage time simulations for random networks
Figure 6: Example of a randomly generated graph with $N=30$ vertices based on the procedure introduced in [67] to obtain networks with clear starting and end points. Both the network topology and the rates are obtained by random processes.
To practically demonstrate how first-passage time distributions are related to the eigenvalue structure of the generator matrix, we now consider random networks of increasing size $N$. We will consider both forward and backward biases as well as both reversible and irreversible dynamics. This requires networks with a source at vertex $1$ and a sink at vertex $N$ that are relatively far apart. Since random graphs generated by Erdős–Rényi-like models tend to give small graph distances in connected graphs, we employ the graph-generation procedure introduced in [67], which yields random graphs such as the one shown in Fig. 6.
In the reversible case, all transitions are bidirectional and rates satisfy detailed balance. In the non-reversible case, directed edges are assigned independently, and backward transitions are included with a fixed probability $p$, so that detailed balance is generally broken. The rates for the edges $(i,j)$ are then assigned as:
$k_{ij}=k_{ij}^{0}e^{b\frac{\text{d}(i,N)-\text{d}(j,N)}{2}},$ (54)
where $d(i,N)$ and $d(j,N)$ denotes the graph distance from vertex $j$ and $i$ to the absorbing target $N$, respectively, and $b$ controls the bias towards ($b>0$) or away from ($b<0$) the target. The prefactors $k_{ij}^{0}$ are drawn from a uniform distribution,
$k_{ij}^{0}\sim U([1-\alpha,1+\alpha]),$ (55)
where $\alpha<1$ sets the variability of the rates. In the reversible case, the prefactors are chosen symmetrically, $k_{ij}^{0}=k_{ji}^{0}$, while in the general case, they are drawn independently.
The stochastic dynamics are simulated using the Gillespie algorithm [78] and the eigenvalues of the corresponding Laplace matrices are computed numerically. Although the network structure and the rates are randomly generated, the resulting matrix retains the Laplacian structure, so its entries are not independent. Hence, simple limit laws for the eigenvalues of random matrices do not apply. More details on the simulations are given in Supplement VI. The code and the data are publicly available [92].
Our results are shown in Fig. 7. The histograms display first-passage time distributions obtained from $10^{5}$ trajectories on representative networks with increasing vertex number $N$. The box plots show the dominance ratio of the principal eigenvalue,
$R=\frac{\lambda_{1}^{-1}}{\sum_{i}\lambda_{i}^{-1}},$ (56)
computed for $100$ independent network realizations for each $N$. The inset diagrams show the corresponding coefficients of variation,
$\mathrm{CV}_{\tau}=\frac{\sigma_{\tau}}{\langle\tau\rangle},$ (57)
obtained from $10^{4}$ first-passage time samples for each network.
The results for reversible networks of increasing size with a forward bias in the rates ($b=1$) are shown in Fig. 7(a): as expected, the FPT distributions converge to the deterministic limit (histograms). The eigenvalue ratio (box plot) and the coefficient of variation of the first-passage time (inset) both vanish with increasing $N$. Fig. 7(b) shows the limiting behavoir for reversible networks with a backward bias b=1b=-1: the exponential limit arises already for moderate system sizes (histograms). Accordingly, the principal eigenvalue dominates the spectrum reflected by the eigenvalue ratio (box plots) and the coefficient of variation (inset) approaching unity.
Fig 7(c) shows that for non-reversible networks, in which reverse transitions are present with probability $p=0.5$, the results for a forward bias in the rates ($b=1$) are qualitatively identical to the reversible case. Again, the deterministic limit arises (histograms) with vanishing eigenvalue ratio (histogram) and coefficient of variation (inset). However, as shown in Fig. 7(d), for non-reversible networks with a backward bias (b=1b=-1), the limiting first-passage time distributions do not generically approach the exponential limit (histograms) and the eigenvalue ratio and the coefficient of variation do not converge to $1$. The reason for this is that unidirectional transitions at bottleneck positions in the network can lead to metastable clusters and checkpoint-like transitions between them. Hence, it becomes transparent that this case has to be handled with care and that depending on non-reversible transitions at crucial points, one should expect a non-generic limit.
Figure 7: First-passage times and eigenvalue ratios for random networks of increasing size. (a) For reversible networks with forward bias, the FPT statistics approaches the deterministic limit as system size NN increases. This is reflected by a vanishing eigenvalue ratio (box plot) and a vanishing coefficient of variation (inset). (b) For reversible networks with backward bias, the FPT statistics approaches the exponential limit as system size NN increases, reflected in a dominant principal eigenvalue. The coefficient of variation approaches 11. (c) For non-reversible networks with forward bias, the FPT statistics approaches the deterministic limit as system size NN increases, similar to case (a). (d) For non-reversible networks with backward bias, the FPT limit is not generic. The eigenvalue ratio neither approaches 0 nor 11. The histograms show the FPT distribution for $10^{5}$ runs on a single network and the boxplots are obtained from 100100 networks on which $10^{4}$ first-passage times were generated.
IV Discussion
It has been reported before for kinetic proofreading networks that two simple limiting distributions arise in the limit of large networks [40, 41]. This observation is surprising in the light of the immense structural complexity of biochemical reaction networks [93]. It calls for a more thorough mathematical explanation, since FPTs in Markovian networks correspond to so-called phase-type distributions, which are dense in the class of distributions with non-negative support [80]. Consequently, they can, a priori, approximate any possible shape of an FPT distribution arbitrarily well.
Here we have established that these limits originate from the structure of the eigenspectrum of the generator matrix. We introduced a condition on the ratio of the weights of two-tree spanning forests, the macroscopic forest condition (Sec. III.2), which allowed us to relate the FPT cumulants to sums over the eigenvalues in the limit $N\to\infty$. From this relationship, we deduced that if infinitely many eigenvalues contribute non-vanishingly to the sum, the FPT distribution converges to the deterministic limit (Sec. III.3).
For strongly connected networks with locally finite outgoing rates, a finite conductance is sufficient to ensure convergence to this deterministic regime. For networks that are not strongly connected, the analysis must be performed separately for each strongly connected component. If the FPTs across components vanish compared to the global FPT as $N\to\infty$, the total FPT distribution becomes a sum of delta peaks corresponding to the distinct traversal paths through these components. As a biological example, embryonic development may be viewed as a system undergoing many irreversible transitions. Although the mean completion time is sensitive to environmental factors such as temperature, the distribution of completion times remains narrowly concentrated around it [94, 95, 96, 97, 98].
The exponential limit, on the other hand, arises when the generator matrix has a single dominant eigenvalue (Sec. III.4). We demonstrated that, if the macroscopic network condition holds, showing that the coefficient of variation approaches unity is sufficient to guarantee convergence to an exponential distribution. Moreover, we established that, at least for reversible networks (i.e., those obeying detailed balance in the steady state), the intuitive interpretation of a backward bias as the mechanism leading to exponential behavior can be formulated as a precise mathematical condition involving the ratio of the MFPTs in the forward and reverse directions.
The physical expectation that a forward bias generically leads to the deterministic limit [40, 41] is mathematically subtle, as shown in Sec. III.5. Instead, one can consider the mean residual life of the process: it long-time limit corresponds to the principal eigenvalue. So if the mean residual life vanishes for large times compared to the mean first-passage time, one can conclude that the eigenvalue ratio must also vanish. The result yields a formal condition for the deterministic limit in terms of a forward bias. While conceptually clear, this formulation is not directly tied to structural network properties, for instance the statistics of the transition rates.
Simulations of first-passage times in random networks with random rates confirm that the deterministic (delta-like) limit emerges robustly in networks with a forward bias. In these cases, the ratio of the principal eigenvalue contribution to the total spectral weight, vanishes with increasing system size, in agreement with the spectral analysis. Conversely, in reversible networks a backward bias generically leads to dominance of a single eigenvalue and hence to exponential first-passage statistics, again as predicted by the spectral analysis.
In irreversible networks, however, the global structure of the graph can qualitatively alter this picture. A backward bias may generate metastable clusters connected by effectively irreversible transitions. These irreversible steps act as stochastic checkpoints, segmenting the dynamics into successive stages and thereby preventing generic spectral concentration. As a consequence, deviations from the simple exponential limit can arise despite the presence of a backward bias. A systematic classification of the backward-biased irreversible regime, in particular the role of metastable clustering and effective dynamical checkpoints, represents an interesting direction for future work.
Overall, we conclude that the universality of these limiting first-passage time distributions holds under clearly identifiable structural conditions. First, the large-network limit can yield universal behavior only if the process starts sufficiently far away from the target state. Otherwise, local non-macroscopic structures dominate. Thus, a condition such as the macroscopic forest condition is needed for universality. Second, irreversible transitions can break this universality: while a network composed of strongly connected microscopic subnetworks yields several delta peaks, if multiple macroscopic strongly connected components exist with irreversible connections between them, the resulting FPT distribution becomes non-generic. One of the simplest example would be a system with two macroscopic components, each individually leading to exponential FPTs, but connected by an irreversible transition. The total FPT distribution in the limit $N\to\infty$ then becomes a convolution of two exponentials — neither purely exponential nor deterministic. Finally, universality cannot be expected if arbitrarily large outgoing rates are allowed (compare the example in Supplement V).
Understanding how simple macroscopic laws emerge from the intricate dynamics of microscopic constituents lies at the heart of statistical physics. While this question was originally treated within the context of equilibrium systems, recent decades have seen remarkable progress in extending these principles to non-equilibrium settings [57, 99, 100, 58], such as living systems [101, 102]. The methods established here might be helpful also in the context of stochastic thermodynamics, because they are valid for arbitrary Markov networks.
Our results highlight that even in high-dimensional Markovian networks, the interplay between spectral properties of the generator and the underlying graph structure governs the emergence of simple limiting first-passage times. For large and complex networks, approaches based on spectral theory [59, 90] and complementary graph-theoretical formalisms [69, 67] therefore provide a powerful unified formalism for uncovering general mechanisms that so far remained hidden in the enormous microscopic parameter space. These insights set the stage for extending the present results to broader classes of stochastic systems and for exploring universal behavior beyond the regimes addressed here.
Acknowledgements.
JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support. We also acknowledge support by the Max Planck School Matter to Life funded by the Dieter Schwarz Foundation and the Max Planck Society.